Determine the mass of 68.1% Na3PO4 solution that contains 9.26 g of Na+.

Answer: _____g solution
Answer #1

Molar mass of Na = 22.99 g/mol

mass(Na)= 9.26 g

use:

number of mol of Na,

n = mass of Na/molar mass of Na

=(9.26 g)/(22.99 g/mol)

= 0.4028 mol

This is number of moles of Na

one mole of Na3PO4 contains 3 mole of Na

use:

number of moles of Na3PO4 = number of moles of Na3PO4 / 3

= 0.4028 / 3

= 0.1343

Molar mass of Na3PO4,

MM = 3*MM(Na) + 1*MM(P) + 4*MM(O)

= 3*22.99 + 1*30.97 + 4*16.0

= 163.94 g/mol

use:

mass of Na3PO4,

m = number of mol * molar mass

= 0.1343 mol * 1.639*10^2 g/mol

= 22.01 g

Now use:

Mass % of Na3PO4 = mass of Na3PO4 * 100 / mass of solution

68.1 = 22.01 * 100 / mass of solution

mass of solution = 32.3 g

Answer: 32.3 g